Ask your own question, for FREE!
Mathematics 9 Online
OpenStudy (shamil98):

similar question, just want to confirm my answer: The probability that a student graduating from Suburban State University has student loans to pay off after graduation is .60. If two students are randomly selected from this university, what is the probability that neither of them has student loans to pay off after graduation? 4/10 * 3/9 = 0.133 since the total # of students isn't given, just a percentage

OpenStudy (zarkon):

since the population size is supposedly larger sampling with or without replacement will almost be the same. so treat each student as having a 40% chance of not having a loan to pay then compute .4*.4=.16

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
2 days ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
3 days ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!